OpenStudy (anonymous):

Evaluate 6x2 − 2x + 14 when x = -2. A. -8 B. 38 C. 42 D. 160 Help me please :c

OpenStudy (anonymous):

All you have to do is plug in the x value into all of the x's...right?

OpenStudy (anonymous):

yes i think

OpenStudy (anonymous):

6(-2)^2 - 2(-2) +14

OpenStudy (anonymous):

\[\Large 6(~)^2-2(~)+14\] All you have to do is plug in the -2

OpenStudy (anonymous):

Just solve that.

OpenStudy (anonymous):

just replace x=-8 i this eq. as 6(-8)^2-2(-8)+14=6(64)+16+14=384+30=414 similarly replace x=-2 and solve

OpenStudy (anonymous):

its 42 then right?

OpenStudy (anonymous):

your answer will be 42.

OpenStudy (anonymous):

thanks